what is the intercept slope of y+4=3(3x+3)
Answer by lwsshak3(11628)

what is the intercept slope of y+4=3(3x+3)
***
Rewrite equation in to standard slope intercept form: y=mx+b, m=slope, b=y-intercept
..
y+4=3(3x+3)
y+4=9x+9
y=9x+5
slope,m=9
y-intercept,b=5